OpenStudy (anonymous):

use prime factorization to reduce 48/72

OpenStudy (slaaibak):

\[{16 * 3}\over{8* 9}\] \[{2^4 * 3} \over { 2^3 * 3^2}\] \[2 \over 3\]

OpenStudy (anonymous):

thats a little confussing

OpenStudy (slaaibak):

What I did: Take the numerator and denominator individually and use prime factorization. 48: 48/2 = 24 24/2 = 12 12/2 = 6 6/2 = 3 We can not divide by 2, but we can divide by 3: 3/3 = 1 That's as far as we go. So I divided by 2 four times and once by 3: Therefore: 48 = 2 * 2 * 2 * 2 * 3 Now we take 72: 72 / 2 = 36 36 / 2 = 18 18 / 2 = 9 Nine is not divisible by 2, but it is by 3, so: 9 / 3 = 3 3 / 3 = 1 That's as far as we can go. I divided by 2 three times and by 3 two times. Therefore 72 = 2 * 2 * 2 * 3 * 3 So, taking the bigger picture: \[{48 \over 72} = { {2 * 2 * 2 * 2 * 3} \over {2 * 2 * 2 * 3 * 3} }\] Now three two's cancel out and one three cancels out and we get \[2 \over 3\] So, we divided by

OpenStudy (slaaibak):

Sorry, ignore the last sentence, don't know how that got there.

OpenStudy (anonymous):

oh okay